calculate the limit of resolution for the yellow light wavelength 550 nm and blue light wavelength 450 nm. which light produces superior resolution?
Answer #1

Limit of resolution (LR) = 0.61 (wavelength) / numerical aperture

Numerical aperture is not given. It will be fixed for both yellow light and blue light. So we will take it as N.

For yellow light,

LR = 0.61 × 550 / N = 335.5 / N nm

For blue light,

LR = 0.61 × 450 / N = 274.5 /N nm

We can see that limit of resolution of blue light is less than yellow light. It means the superior resolution is that of blue light. Smaller the value of limit of resolution, greater is the resolving power.

Yellow light can resolve two objects placed at a distance of 335.5 nm as two separate objects whereas blue light can resolve two objects placed at a distance of 274.5 nm as two separate objects.

Please rate.